# SeBS-Flow: Benchmarking Serverless Cloud Function Workflows

### v2

@@ Abstract
I Introduction
@@
Serverless computing gained major adoption in the industry [1, 2], with 50-70% of cloud customers using serverless functions and containers [3]. In the Function-as-a-Service (FaaS) programming model, developers implement stateless functions and invoke them through a REST interface. The actual function deployment and resource scheduling becomes the responsibility of the cloud operator: Developers are no longer concerned with managing their applications and are charged only for resources used to handle function invocations. While the primitiveness of FaaS can be an important benefit [1], it is also a major drawback: a single function is insufficient to cover all use cases. Functions must be composed to build larger applications, keep the design modular, or use pre-defined and standardized functions, e.g., for machine learning inference.
@@
Serverless workflows introduced allow to chain and aggregate multiple functions into a single application by creating a graph of functions and automating the execution of a sequence through control and data dependencies. They include control-flow components - conditions and loops - which allows them to represent full computations such as multi-stage machine learning pipelines. Developers implement functions and define the workflow structure in a cloud-specific format. Cloud operators then control the workflow invocation and orchestration, retaining the ability to optimize resource consumption, e.g., through optimized function placement, oversubscription, targeting idle resources, and co-locating functions that depend on each other [4, 5, 6].
@@
Workflows have been adopted by the most popular commercial cloud platforms [7, 8, 9] and make up almost a third of serverless applications [10]. However, just like every FaaS platform is different [11], serverless workflows are quite distinct from each other. Not only the different APIs and incompatible graph syntax and format complicate the software development process, but also fundamentally different programming models: workflow platforms diverge in the statelessness of functions and the static nature of graph definition (Section II-A). Even though FaaS platforms might seem like the same product, they offer drastically different performance, reliability, and cost [11, 12, 13]. With workflows built as an orchestration of functions, their functionality and performance is affected by both orchestration service and existing differences in the underlying compute infrastructure. As such details are hidden, an information gap between developers and providers arises [14]. Thus, the software developers need to conduct extensive performance testing of the cloud services to estimate the performance of their workloads and understand platform limitations up-front, as choosing a certain platform implies significant lock-in [15], with only limited support for testing [2].
@@
TABLE I: Analysis of 72 research papers on serverless workflows with benchmarks.
We propose the first serverless workflows benchmarking suite to support software developers and the quickly growing research activity in serverless workflows. Our work provides a baseline and benchmarking methodology for evaluating and comparing the performance of workflows on different platforms, highlighting their strengths and weaknesses. We examined 72 different research contributions to determine the similarity of their evaluation baselines (Table I). We found that publications use different applications to benchmark the performance of new ideas, do not cover the same classes of workloads, and do not always compare against the same subset of platforms. Without a consistent baseline, comparing research results and establishing the most promising ideas becomes impossible [16]. Benchmarking suites and systems have been proposed for FaaS [11, 12, 17, 18], but a benchmarking suite for serverless workflows has remained an open problem. A comprehensive, consistent, platform-independent, and portable benchmarking suite will support the ongoing research work [19, 16] and enable developers to differentiate between alternative solutions. We establish a unified and portable workflow model to abstract away the differences between different platforms (Section III). We design the benchmarking suite (Section IV) and include six workflow benchmarks based on solutions common in research and industry (Section V). Applications are implemented in our unified workflow model, providing an identical benchmark structure for each platform. We evaluate expressiveness and overhead of our model (Section VI) and use our benchmarking suite to comprehensively evaluate the three major cloud workflow services (Section VII). We follow the FAIR principle [20] and release our benchmark suite on an open-source license, enabling automatic repetition of our experiments, allowing reproducible results, and measuring performance changes in clouds over time. We make the following contributions:
@@
II Background
Serverless workflows introduce multiple new challenges to the software development process due to differences in the workflows platforms (Section II-A). To model workflows, we use the formalism and semantics of Petri Nets (Section II-B).
II-A Developing Serverless Workflows
@@
TABLE II: Key features of serverless workflows platforms.
While software engineers are increasingly interested in serverless applications [21], they encounter a wide range of challenges while developing them, with the first questions about the different capabilities of the platforms arising before starting the implementation [21, 22]: Workflows have been adopted by all major cloud providers, but their implementations are significantly different in capabilities (Table II). We focus on AWS Step Functions, Google Cloud Workflows, and Azure Durable Functions, as they play a leading role.
@@
Figure 1: Workflow invoking function process in parallel, with inputs from zero to three and results written to res.
The most important change is the programming model, affecting the implementation of the workflows, with unknown implications to workflow performance, an important property for developers [21]. As the different implementations are all provider-specific, moving workflows from one platform to another is complicated, causing vendor lock-in [22]. Azure uses the programming model of Durable Functions [23], where the workflow definition is encoded within a regular program structure of an orchestrator. The graph of functions is expressed using a mainstream programming language such as Python, as seen in the example of mapping the elements of input values array to invocations of the process function (Figure 1(a)). The computation model is built on top of stateless activity and stateful entity functions. On the other hand, developers need to define their workflow using a state machine on Google Cloud Workflows and AWS Step Functions. The workflow consists of states representing computations and transitions connecting them. The main states include function invocations, while supplementary states encode control flow. State languages defined with a syntax based on JSON and YAML files can be limited, verbose, and consequently difficult to debug, with missing tool support for testing and debugging already being a problem for developers [24, 2]. The example implementations in Figure 1 demonstrate how simple code snippets can become much more verbose when compared to a native implementation of orchestrator. In Durable Functions, implementing the same behavior requires less work and the single-source implementation is more readable and easier to debug. However, the static form of a state machine gives the cloud provider deep knowledge of the functions executed and their order, allowing for optimizations.
The programming model also has an impact on the billing system. In addition to the cost of executing functions within a workflow, cloud providers charge users for workflow orchestration. In Azure, users have to pay for the duration of the orchestration function. In AWS and Google Cloud, users are charged per each transition of the state machine. Table III shows an overview. Note that we have to estimate the orchestration cost on Azure as billing is at the granularity of complete workflows only.
@@
TABLE III: Pricing according to vendors’ documentation [25, 26, 27, 28, 29]. Orchestration per 1000 transitions.
II-B Workflow Nets
We base our model on workflow nets with data (WFD-nets) [30]. They are an extension of Petri nets, usually used for business workflows. Basing the model on Petri Nets is only one possibility among alternatives such as state machines. We opt for Petri Nets due to their advantages as modeling formalism, such as their graphical nature, formal semantics, and analysis defined. Petri nets [31] describe the flow of information and control in concurrent and asynchronous systems. A Petri net is a triple $T=\langle P,T,F\rangle$ consisting of places P, a finite set of transitions T, and a set of arcs $F\subseteq(P\times T)\cup(T\times P)$. It is a workflow net iff there is a single source place start without incoming arcs, a single sink place without outgoing arcs, and every node is on a path from source to sink [30]. WFD-nets [30] are a tuple $\langle P,T,F,D,r,w,d,grd\rangle$, consisting of a Petri Net $N=\langle P,T,F\rangle$ and additionally containing a set $D$ of data elements on top as well as read, write, and destroy operations on these data elements. Moreover, the guarding function $grd:T\rightarrow G_{D}$ can assign guards to transitions. We show an example in Figure 2 where $t_{1}$ writes data to $x$, while $t_{2}$ and $t_{3}$ read from x. Dynamic system properties are modeled using tokens that are routed through the net. A transition is enabled if tokens are in all its input places $\bullet t=\{p|(p,t)\in F\}$. When it fires, it removes the token(s) from its input place(s) and routes them to its output place(s) $t\bullet=\{p|(t,p)\in F\}$. In our example, $t_{1}$ will be enabled if there is a token in $start$ and put tokens to $p_{1}$ and $p_{2}$, which will enable $t_{2}$ and $t_{3}$.
@@ II-B Workflow Nets
Figure 2: WFD-net with transitions $T=\{t_{1},t_{2},t_{3},t_{4}\}$ and places $P=\{p_{1},p_{2},p_{3},p_{4},start,end\}$
III Serverless Workflows Model
We define a model for serverless workflows that allows developers to implement and analyze a workflow application independent of the platform it will run on, alleviating provider lock-in. The model should encode the control flow and task parallelism, and clearly display the flow of data between functions, aiding developers in detecting scalability bottlenecks and errors, e.g., inconsistent or missing data. Therefore, we define our model on top of WFD-nets [30] (cf. Section II-B) and extend them to be able to express the orchestration by the platform and how data is passed between functions.
Figure 3: Workflow using our model based on WFD-nets.
III-A Transitions
@@
A function transition $sf\in SF$ represents the execution of a serverless function. All function transitions that can run in parallel without any precedence dependencies and their immediate predecessor and successor places make up a workflow phase. There are different possible token routing constructs within one phase of the workflow: A task phase is a sequential routing, consisting of one function transition only. For parallel routing, there are two alternatives: First, a parallel phase can consist of any number of sub-phases that will be executed concurrently. Second, the map phase. Similar to the parallel phase, it can consist of any number of sub-phases, but each sub-phase is executed concurrently on different elements of an input array. Figure 3 shows an example: The map functions compute $y_{i}=\texttt{map}\left(x_{i}\right)$ simultaneously for all $i$. A switch phase uses conditional routing based on values of data by annotating guarding functions to transitions.
@@ III-B Resource Annotations
IV Workflows Benchmark Suite
We now present the design and implementation of SeBS-Flow111 An extended definition and discussion of benchmarks can be found in the Master thesis [32]. . To enable reliable and fair comparison of various workflow platforms, we need to execute the same benchmark implementation on many platforms. However, the platforms exhibit vast differences in the programming model and API of their workflow services (Section II-A). Thus, we define a platform-agnostic workflow definition (Section IV-A) based on our workflow model (Section III). Then, we propose platform-specific generators that transcribe workflows to the respective proprietary definition of the desired platform (Section IV-B). We add the workflow representation and implementation to a serverless benchmark suite (Section IV-C).
IV-A Platform-Agnostic Workflow Definition
Our workflow model encodes the application as Petri Net (cf. Section III). Every phase has a type, relating to one of the available routing constructs (cf. Section III-A). Coordinator transitions encode the order of phases, represented by the next field of phases that describes the consecutive step in the workflow. The workflow will terminate if the next field is not set. Each phase receives the output payload of the previous function as input. We encode the different phases as follows:
@@ IV-A Platform-Agnostic Workflow Definition
[obeytabs=true,tabsize=2,frame=lines,framesep=2mm,fontsize=]jsonres/implementation/map-long.json
(a) Map Statement.
@@
[obeytabs=true,tabsize=2,frame=lines,framesep=2mm,fontsize=]jsonres/implementation/repeat-long.json
(b) Repeat Statement.
@@
[obeytabs=true,tabsize=2,frame=lines,framesep=2mm,fontsize=]jsonres/implementation/switch-long.json
(c) Switch Statement.
Figure 5: Workflow definition language: a portable specification of control-flow and data dependencies.
Task. A task executes a single serverless function.
Map. The map phase concurrently executes the given states one after another on each element of the given array and returns an array again. The phase can define common_parameters from the running variable that will be passed in addition to the array element. Listing 5(a) shows an example with the process_names phase: for each element of customers, the function short is executed concurrently. Only after all functions have terminated, the coordinator will transition to the next phase, which in this case is list_emails.
@@
Repeat. A repeat phase executes a function a given number of times. This syntactic sugar eases the modeling of a chain of tasks. Listing 5(b) presents an example where the function process is invoked 10 times, and the return payload of the $i^{th}$ invocation is passed onto the $i$+$1^{th}$ execution.
Switch. The switch phase dynamically decides the next phase at runtime depending on the given condition. Listing 5(c) presents a simple switch phase where different functions are executed depending on the running variable data.length.
Parallel. This higher-level phase executes sub-workflows, consisting of any of the phases, concurrently.
IV-B Platform-Specific Transcription
We map the six phases building a serverless workflow to different features of the modeling language on each platform.
IV-B1 AWS
The most notable difficulty when transcribing our definition to the state machine definition of AWS Step Functions is the loop phase. Step Functions do not inherently support sequential array iteration. Their official documentation suggests using an additional serverless function that iterates over a given range [33], which is inefficient. Thus, we use the AWS map state and configure it to traverse the given array sequentially, yielding the semantics of a loop. A downside of this approach is that the input to each function is the same, i.e., consecutively executed functions can observe the results of computations of their predecessors only if uploaded to the object storage.
IV-B2 Google Cloud
Google Cloud Workflows do not natively support a task type. Instead, the recommended approach for invoking Cloud Functions [34] is to create a state performing a POST request and providing the trigger URL of the desired function as input. However, this requires additional states for each task and map to parse the HTTP response of a function and assign results. Moreover, the parallel map execution accepts only other workflows and not states, which requires creating another sub-workflow, even if it contains only a single function to be invoked. Finally, there is no mechanism for passing additional arguments to a map function, which is necessary for us to track measurements. As a workaround, the input array is zipped together with an array consisting of the additional parameter passed by the benchmarking infrastructure.
IV-B3 Azure

### v3

# {textblock}

@@ Abstract.
benchmark, serverless, function-as-a-service, faas, workflow, orchestration, serverless DAG
copyright: nonejournalyear: 2025copyright: ccconference: Twentieth European Conference on Computer Systems; March 30-April 3, 2025; Rotterdam, Netherlandsbooktitle: Twentieth European Conference on Computer Systems (EuroSys ’25), March 30-April 3, 2025, Rotterdam, Netherlandsdoi: 10.1145/3689031.3717465isbn: 979-8-4007-1196-1/2025/03ccs: Computer systems organization Cloud computingccs: Networks Cloud computingccs: Software and its engineering Software performanceccs: General and reference Performanceccs: General and reference Metricsccs: General and reference Measurementccs: General and reference Evaluation
1. Introduction
@@
Serverless computing gained major adoption in the industry (Eismann et al., 2021; Leitner et al., 2019), with 50-70% of cloud customers using serverless functions and containers (Datadog, 2024). In the Function-as-a-Service (FaaS) programming model, developers implement stateless functions and invoke them through a REST interface. The actual function deployment and resource scheduling becomes the responsibility of the cloud operator: Developers are no longer concerned with managing their applications and are charged only for resources used to handle function invocations. While the primitiveness of FaaS can be an important benefit (Eismann et al., 2021), it is also a major drawback: a single function is insufficient to cover all use cases. Functions must be composed to build larger applications, keep the design modular, or use pre-defined and standardized functions, e.g., for machine learning inference.
@@
Serverless workflows allow to chain and aggregate multiple functions into a single application by creating a graph of functions and automating the execution of a sequence through control and data dependencies. They include control-flow components - conditions and loops - which allows them to represent full computations such as multi-stage machine learning pipelines. Developers implement functions and define the workflow structure in a cloud-specific format. Cloud operators then control the workflow invocation and orchestration, retaining the ability to optimize resource consumption, e.g., through optimized function placement, oversubscription, targeting idle resources, and co-locating functions that depend on each other (Mahgoub et al., 2021; Akkus et al., 2018; Copik et al., 2024).
@@
Workflows have been adopted by the most popular commercial cloud platforms (aws, 2016; azu, 2019; gcp, 2020) and make up almost a third of serverless applications (Eismann et al., 2022). However, just like every FaaS platform is different (Copik et al., 2021), serverless workflows are quite distinct from each other. Not only the different APIs and incompatible graph syntax and format complicate the software development process, but also fundamentally different programming models: workflow platforms diverge in the statelessness of functions and the static nature of graph definition (Section 2.1). Even though FaaS platforms might seem like the same product, they offer drastically different performance, reliability, and cost (Copik et al., 2021; Maissen et al., 2020; Wang et al., 2018). With workflows built as an orchestration of functions, their functionality and performance is affected by both orchestration service and existing differences in the underlying compute infrastructure. As such details are hidden, an information gap between developers and providers arises (Wen et al., 2023). Thus, the software developers need to conduct extensive performance testing of the cloud services to estimate the performance of their workloads and understand platform limitations up-front, as choosing a certain platform implies significant lock-in (Adzic and Chatley, 2017), with only limited support for testing (Leitner et al., 2019).
@@
Table 1. Analysis of 72 research papers on serverless workflows with benchmarks.
We propose the first serverless workflows benchmarking suite to support software developers and the quickly growing research activity in serverless workflows. Our work provides a baseline and benchmarking methodology for evaluating and comparing the performance of workflows on different platforms, highlighting their strengths and weaknesses. We examined 72 different research contributions to determine the similarity of their evaluation baselines (Table 1). We found that publications use different applications to benchmark the performance of new ideas, do not cover the same classes of workloads, and do not always compare against the same subset of platforms. Without a consistent baseline, comparing research results and establishing the most promising ideas becomes impossible (Scheuner and Leitner, 2020). Benchmarking suites and systems have been proposed for FaaS (Copik et al., 2021; Maissen et al., 2020; Somu et al., 2020; Kim and Lee, 2019), but a benchmarking suite for serverless workflows has remained an open problem. A comprehensive, consistent, platform-independent, and portable benchmarking suite will support the ongoing research work (Papadopoulos et al., 2021; Scheuner and Leitner, 2020) and enable developers to differentiate between alternative solutions. We establish a unified and portable workflow model to abstract away the differences between different platforms (Section 3). We design the benchmarking suite (Section 4) and include six workflow benchmarks based on solutions common in research and industry (Section 5). Applications are implemented in our unified workflow model, providing an identical benchmark structure for each platform. We evaluate expressiveness and overhead of our model (Section 6) and use our benchmarking suite to comprehensively evaluate the three major cloud workflow services (Section 7). We follow the FAIR principle (Wilkinson et al., 2016) and release our benchmark suite on an open-source license, enabling automatic repetition of our experiments, allowing reproducible results, and measuring performance changes in clouds over time. We make the following contributions:
@@
2. Background
Serverless workflows introduce multiple new challenges to the software development process due to differences in the workflows platforms (Section 2.1). To model workflows, we use the formalism and semantics of Petri Nets (Section 2.2).
2.1. Developing Serverless Workflows
@@
Table 2. Key features of serverless workflows platforms.
While software engineers are increasingly interested in serverless applications (Wen et al., 2021), they encounter a wide range of challenges while developing them, with the first questions about the different capabilities of the platforms arising before starting the implementation (Wen et al., 2021; Sampe et al., 2021): Workflows have been adopted by all major cloud providers, but their implementations are significantly different in capabilities (Table 2). We focus on AWS Step Functions, Google Cloud Workflows, and Azure Durable Functions, as they play a leading role.
@@
Figure 1. Workflow invoking function process in parallel, with inputs from zero to three and results written to res.
The most important change is the programming model, affecting the implementation of the workflows, with unknown implications to workflow performance, an important property for developers (Wen et al., 2021). As the different implementations are all provider-specific, moving workflows from one platform to another is complicated, causing vendor lock-in (Sampe et al., 2021). Azure uses the programming model of Durable Functions (Burckhardt et al., 2021), where the workflow definition is encoded within a regular program structure of an orchestrator. The graph of functions is expressed using a mainstream programming language such as Python, as seen in the example of mapping the elements of input values array to invocations of the process function (Figure 1(a)). The computation model is built on top of stateless activity and stateful entity functions. On the other hand, developers need to define their workflow using a state machine on Google Cloud Workflows and AWS Step Functions. The workflow consists of states representing computations and transitions connecting them. The main states include function invocations, while supplementary states encode control flow. State languages defined with a syntax based on JSON and YAML files can be limited, verbose, and consequently difficult to debug, with missing tool support for testing and debugging already being a problem for developers (Wen et al., 2022; Leitner et al., 2019). The example implementations in Figure 1 demonstrate how simple code snippets can become much more verbose when compared to a native implementation of orchestrator. In Durable Functions, implementing the same behavior requires less work and the single-source implementation is more readable and easier to debug. However, the static form of a state machine gives the cloud provider deep knowledge of the functions executed and their order, allowing for optimizations.
The programming model also has an impact on the billing system. In addition to the cost of executing functions within a workflow, cloud providers charge users for workflow orchestration. In Azure, users have to pay for the duration of the orchestration function. In AWS and Google Cloud, users are charged per each transition of the state machine. Table 3 shows an overview. Note that we have to estimate the orchestration cost on Azure as billing is at the granularity of complete workflows only.
@@
Table 3. Pricing according to vendors’ documentation (AWS Step Functions Pricing, 2024; Azure Functions Pricing, 2024; Google Cloud Workflows Pricing, 2024; Google Cloud, 2024a; AWS Lambda Pricing, 2024). Orchestration per 1000 transitions.
2.2. Workflow Nets
We base our model on workflow nets with data (WFD-nets) (Trčka et al., 2009). They are an extension of Petri nets, usually used for business workflows. Basing the model on Petri Nets is only one possibility among alternatives such as state machines. We opt for Petri Nets due to their advantages as modeling formalism, such as their graphical nature, formal semantics, and analysis defined. Petri nets (Peterson, 1977) describe the flow of information and control in concurrent and asynchronous systems. A Petri net is a triple $T=\langle P,T,F\rangle$ consisting of places P, a finite set of transitions T, and a set of arcs $F\subseteq(P\times T)\cup(T\times P)$. It is a workflow net iff there is a single source place start without incoming arcs, a single sink place without outgoing arcs, and every node is on a path from source to sink (Trcka et al., 2008). WFD-nets (Trcka et al., 2008) are a tuple $\langle P,T,F,D,r,w,d,grd\rangle$, consisting of a Petri Net $N=\langle P,T,F\rangle$ and additionally containing a set $D$ of data elements on top as well as read, write, and destroy operations on these data elements. Moreover, the guarding function $grd:T\rightarrow G_{D}$ can assign guards to transitions. We show an example in Figure 2 where $t_{1}$ writes data to $x$, while $t_{2}$ and $t_{3}$ read from x. Dynamic system properties are modeled using tokens that are routed through the net. A transition is enabled if tokens are in all its input places $\bullet t=\{p|(p,t)\in F\}$. When it fires, it removes the token(s) from its input place(s) and routes them to its output place(s) $t\bullet=\{p|(t,p)\in F\}$. In our example, $t_{1}$ will be enabled if there is a token in $start$ and put tokens to $p_{1}$ and $p_{2}$, which will enable $t_{2}$ and $t_{3}$.
@@ 2.2. Workflow Nets
Figure 2. WFD-net with transitions $T=\{t_{1},t_{2},t_{3},t_{4}\}$ and places $P=\{p_{1},p_{2},p_{3},p_{4},start,end\}$
3. Serverless Workflows Model
We define a model for serverless workflows that allows developers to implement and analyze a workflow application independent of the platform it will run on, alleviating provider lock-in. The model should encode the control flow and task parallelism, and clearly display the flow of data between functions, aiding developers in detecting scalability bottlenecks and errors, e.g., inconsistent or missing data. Therefore, we define our model on top of WFD-nets (Trcka et al., 2008) (cf. Section 2.2) and extend them to be able to express the orchestration by the platform and how data is passed between functions.
Figure 3. Workflow using our model based on WFD-nets.
3.1. Transitions
@@
A function transition $sf\in SF$ represents the execution of a serverless function. All function transitions that can run in parallel without any precedence dependencies and their immediate predecessor and successor places make up a workflow phase. There are different possible token routing constructs within one phase of the workflow: A task phase is a sequential routing, consisting of one function transition only. For parallel routing, there are two alternatives: First, a parallel phase can consist of any number of sub-phases that will be executed concurrently. Second, the map phase: Similar to the parallel phase, it can consist of any number of sub-phases, but each sub-phase is executed concurrently on different elements of an input array. Figure 3 shows an example: The map functions compute $y_{i}=\texttt{map}\left(x_{i}\right)$ simultaneously for all $i$. A switch phase uses conditional routing based on values of data by annotating guarding functions to transitions.
@@ 3.2. Resource Annotations
  • [Uncaptioned image] NoSQL. Data stored in NoSQL key-value storage provides low-latency data storage.
@@
Formally, we define the set of resource annotations $A=\{o,n,p,t,r\}$ as additional element of the tuple of a WFD-net, with $o$ representing data passing via the object storage, $n$ via NoSQL, $p$ via the invocation payload, $t$ transparently, and $r$ via reference. We define the corresponding resource annotation functions for reading and writing data as $ra$ and $rw$ as follows and also add them to the tuple of a WFD-net:
$\text{ra}:\{(t,d)\in T\times D\mid d\in r(t)\}\to A$
$\text{rw}:\{(t,d)\in T\times D\mid d\in w(t)\}\to A$
This means that each pair of a transition and a data element $(t,d)$, with $d$ being read or written by $t$, respectively, is assigned a resource annotation $a\in A$. By adding resource annotations, we do not change the behavior of the WFD-net. However, we enable checking the consistency of data accesses, for example, if the same data object is written and read using the same resource annotation.
@@
4. Workflows Benchmark Suite
We now present the design and implementation of SeBS-Flow111 An extended definition and discussion of benchmarks can be found in the Master thesis (lau, 2022). . To enable reliable and fair comparison of various workflow platforms, we need to execute the same benchmark implementation on many platforms. However, the platforms exhibit vast differences in the programming model and API of their workflow services (Section 2.1). Thus, we define a platform-agnostic workflow definition (Section 4.1) based on our workflow model (Section 3). Then, we propose platform-specific generators that transcribe workflows to the respective proprietary definition of the desired platform (Section 4.2). We add the workflow representation and implementation to a serverless benchmark suite (Section 4.3).
4.1. Platform-Agnostic Workflow Definition
Our workflow model encodes the application as Petri Net (cf. Section 3). To define workflows in SeBS-Flow conforming to our model, we use a JSON syntax. Every phase has a type, relating to one of the available routing constructs (cf. Section 3.1). Coordinator transitions encode the order of phases, represented by the next field of phases that describes the consecutive step in the workflow. The next field refers to the phase name to be executed after, and the workflow terminates if this field is not set. Each phase receives the output payload of the previous function as input. This means that function implementations need to conform to the resource annotations as defined in the workflow model and download and upload data as needed accordingly. We encode the different phases as follows:
@@ 4.1. Platform-Agnostic Workflow Definition
[obeytabs=true,tabsize=2,frame=lines,framesep=2mm,fontsize=]jsonres/implementation/task.json
(a) Task Statement.
@@
[obeytabs=true,tabsize=2,frame=lines,framesep=2mm,fontsize=]jsonres/implementation/map.json
(b) Map Statement.
@@
[obeytabs=true,tabsize=2,frame=lines,framesep=2mm,fontsize=]jsonres/implementation/complete-workflow.json
(c) Workflow from Figure 3.
Figure 5. Workflow definition language: a portable specification of control-flow and data dependencies.
Task. A task executes a single serverless function, constituting a sequential routing. Listing 5(a) shows an example with the compute_phase executing the function compute.
Map. The map phase is a parallel routing construct and concurrently executes the given states one after another on each element of the given array and returns an array again. The phase can define common_parameters from the running variable that will be passed in addition to the array element. Listing 5(b) shows an example with the process_names phase: for each element of customers, the function short is executed concurrently. Only after all functions have terminated, the coordinator will transition to the next phase, which in this case is list_emails.
@@
Repeat. A repeat phase executes a function a given number of times. This syntactic sugar eases modeling a chain of tasks.
Switch. The switch phase is a conditional routing, deciding the next phase dynamically at runtime based on the given condition. The different cases are evaluated after another, with the first one fulfilling the condition being executed.
Parallel. This higher-level phase corresponds to a parallel routing and executes sub-workflows, consisting of any of the phases, concurrently.
We show an example of a complete workflow definition in Listing 5(c), encoding the same workflow as shown in Figure 3. The root entry specifies the name of the phase that should be executed first, in this case the generate_phase. The states entry then contains all phases of the workflow. As mentioned above, each phase receives the output payload of the previous function as input. Therefore, the data movement is encapsulated in the functions and not controlled by the workflow orchestration. Only in the case of the map phase, we explicitly specify which array is used for distributing its elements to single functions. The level of parallelism is then decided dynamically at runtime depending on the size of the given array.
4.2. Platform-Specific Transcription
We map the six phases building a serverless workflow to different features of the modeling language on each platform. We evaluate the overhead introduced by necessary adaptations to platforms in Section 6.
4.2.1. AWS
The most notable difficulty when transcribing our definition to the state machine definition of AWS Step Functions is the loop phase. Step Functions do not inherently support sequential array iteration. Their official documentation suggests using an additional serverless function that iterates over a given range (AWS, 2024), which is inefficient. Thus, we use the AWS map state and configure it to traverse the given array sequentially, yielding the semantics of a loop. A downside of this approach is that the input to each function is the same, i.e., consecutively executed functions can observe the results of computations of their predecessors only if uploaded to the object storage.
4.2.2. Google Cloud
Google Cloud Workflows do not natively support a task type. Instead, the recommended approach for invoking Cloud Functions (Google Cloud, 2024b) is to create a state performing a POST request and providing the trigger URL of the desired function as input. However, this requires additional states for each task and map to parse the HTTP response of a function and assign results. Moreover, the parallel map execution accepts only other workflows and not states, which requires creating another sub-workflow, even if it contains only a single function to be invoked. Finally, there is no mechanism for passing additional arguments to a map function, which is necessary for us to track measurements. As a workaround, the input array is zipped together with an array consisting of the additional parameter passed by the benchmarking infrastructure.
4.2.3. Azure